### v2

@@ 5.3 Attack Filtering
In addition to the measurement update, the attacks could also be filtered out via the intersections in the agreement protocol (8), which could yield an empty agreement set $\mathcal{I}_{k}^{h}$ in case of inconsistencies between the measurement updates of the sensors in $\mathsf{J}_{h}$. However, unlike the measurement update, the attacks may not be always identified whenever some of the agreement sets are empty. This is because a compromised sensor $i$ could be a part of multiple agreement sets $\mathcal{I}_{k}^{h}$, for $h\in\mathbb{Z}_{[1,n_{\mathsf{J}}]}$, some of which may turn out to be non-empty. One can identify the attack $a_{i}(k)$ if, for every $h\in\mathbb{Z}_{[1,n_{\mathsf{J}}]}$ such that $i\in\mathsf{J}_{h}$, we have $\mathcal{I}_{k}^{h}=\emptyset$.
@@
¯$\bar{\mathcal{I}}_{k}^{h,\mathsf{A}_{k}}=\langle c_{\mathcal{I}}^{h,\mathsf{A}_{k}}(k),G_{\mathcal{I}}^{h,\mathsf{A}_{k}}(k)\rangle.$
@@
¯$\bar{\mathcal{I}}_{k}^{h,\mathsf{S}_{k}}=\langle c_{\mathcal{I}}^{h,\mathsf{S}_{k}}(k),G_{\mathcal{I}}^{h,\mathsf{S}_{k}}(k)\rangle.$
@@
The proof follows from a similar lines of argument as the proof of Theorem 3. ∎
@@
Notice that $c_{\mathcal{I}}^{h,\mathsf{A}_{k}}(k)$, which is the center of the overbounding zonotope ¯$\bar{\mathcal{I}}_{k}^{h,\mathsf{A}_{k}}$, is actually the Chebyshev center of the agreement set $\mathcal{I}_{k}^{h,\mathsf{A}_{k}}$. This agreement set is formed by the intersections of the measurement update sets of the attacked sensors $\mathsf{A}_{k}$. The centers of these measurement update sets are influenced by the corresponding attack signals $a_{j}(k)$. Therefore, to achieve $\mathcal{I}_{k}^{h}\neq\emptyset$, it is necessary that the attacker coordinates the attacks on sensors in such a way that the condition of Proposition 2 is violated.
We remark again that from the attacker’s perspective, the condition of Proposition 2 is necessary for $\mathcal{I}_{k}^{h}\neq\emptyset$, and is not sufficient. This points to several interesting open problems for designing effective attacks. Beyond Theorem 3 and Propositions 1 and 2, our future work will investigate whether a sufficient condition could be derived for the attacker to ensure that both the measurement updates and the agreement sets are non-empty. Moreover, the complexity of generating such attacks is also one of the important questions deferred for our future work.
@@ 5.4 Algorithm for Attack Filtering and Identification
A notable contribution of the set-based state estimation scheme in this paper over other secure schemes is Assumption 1(i), which allows the attacker to compromise not only up to $p-1$ sensors at each time instant but also different subsets of sensors at different times. To this end, we remark that we can detect only those compromised sensors that are injected with non-stealthy attack signals, where by non-stealthiness we mean the attack signals that satisfy the conditions of Theorem 3 and Propositions 1 and 2. The attack identification algorithm is fairly simple and can be summarized as in Algorithm 2.
@@
The estimated safe subset $\hat{\mathsf{S}}_{k}$ contains the true safe subset $\mathsf{S}_{k}$ at time $k$, i.e., $\mathsf{S}_{k}\subseteq\hat{\mathsf{S}}_{k}$. If there are sensors that are injected with small-valued stealthy attack signals, our detection algorithm fails to recognize those attacks and considers those sensors to be uncompromised. Therefore, the estimated set of identified attacked sensors $\hat{\mathsf{A}}_{k}$ is only a subset of the true set of attacked sensors $\mathsf{A}_{k}=\mathbb{Z}_{[1,p]}\setminus\mathsf{S}_{k}$.
@@
In Algorithm 2, we identify a subset of attacked sensors at every time $k$. However, in the time-invariant attack setting where the attacker does not change the subset of sensors to compromise at every time instant $k$, Algorithm 2 can be adapted to cumulatively detect and remove the attacked sensors over time. This can have application in sensor fault detection and isolation as faults can be considered as naive, time-invariant attacks.
6 Stability of the Estimated Set
To assist in the stability analysis of the secure zonotopic state estimation algorithm, in the sense that the estimated set $\hat{\mathcal{X}}_{k}$ containing the true state $x(k)$ is bounded with respect to time $k\in\mathbb{Z}_{\geq 0}$ given that the initial set $\mathcal{X}$ is bounded, we employ a secure point-based state estimator in parallel. Such an estimator provides a point-based estimate of the state and ensures that the bound on the estimation error is unaffected by the attack signal. However, for this section, we restrict Assumption 1(i) by considering $2q<p$. In the following, we define a point-based secure estimator and use it to slightly modify our proposed S3E algorithm, which ensures that the set-based estimate is stable.
6.1 Existence of a Secure Point-based State Estimator
To guarantee stability of S3E algorithm, we employ a point-based estimator which is assumed to be secure with respect to the attack signals $a_{i}$ and robust with respect to disturbances $w$ and noise $v$ as follows.
Assumption 2.
There exists a secure point-based state estimator $E:\mathbb{R}^{m}\times\mathcal{U}\rightarrow\mathbb{R}^{n_{x}}$ which provides a point-based state estimate $\hat{x}(k)=E(y_{1}(k),\dots,y_{p}(k),u(k))$ to system (3) satisfying the following for $k\in\mathbb{Z}_{\geq 0}$,
$\|\hat{x}(k)-x(k)\|\leq\beta(\|\hat{x}(0)-x(0)\|,k)+{\alpha\max(\bar{w},\bar{v})}$ (16)
for some class $\mathcal{KL}$ function $\beta$ and $\alpha\in\mathbb{R}_{\geq 0}$, where $\|w(k)\|_{\infty}\leq\bar{w}$ and $\|v(k)\|_{\infty}\leq\bar{v}$ for all $k\in\mathbb{Z}_{\geq 0}$.
Note that in Assumption 2, the upper bound on the state estimation error is not affected by the attack signals $a_{i}$. We call estimators that possess this property secure and results exist in the literature for linear systems (e.g., [12, 27, 21, 16, 9]). So far, these results require at most $q\leq\lceil p/2\rceil-1$ number of attacked sensors, i.e., strictly less than half the number of sensors can be attacked. Thereby, this imposes a stricter requirement than Assumption 1(i), which we state below.
Assumption 3.
The number of attacked sensors $q\in\mathbb{Z}_{\geq 0}$ satisfies $2q<p$, where $p\in\mathbb{Z}_{\geq 0}$ is the total number of sensors. The integers $p$ and $q$ are known, but the exact attacked sensors are unknown.
In the following subsection, we provide a stability analysis of S3E algorithm by using a secure point-based state estimator in parallel.
6.2 Modified Estimate
We modify the estimate as
$\hat{\mathcal{X}}^{\mathrm{mod}}_{k}=\hat{\mathcal{X}}_{k}\cap\mathcal{H}(\hat{x}(k),\Delta_{k})$ (17)
where $\hat{\mathcal{X}}_{k}$ is the estimate in (9) and $\mathcal{H}(\hat{x}(k),\Delta_{k})$ denotes a hypercube with center $\hat{x}(k)$, which is the estimate provided by a secure point-based estimator under Assumption 2, and radius
$\Delta_{k}\coloneq\beta(\|x(0)-\hat{x}(0)\|,k)+\alpha\max(\bar{w},\bar{v})$ (18)
where $\beta\in\mathcal{KL}$, $\alpha\in\mathbb{R}_{\geq 0}$, $\bar{w}\in\mathbb{R}_{\geq 0}$ and $\bar{v}\in\mathbb{R}_{\geq 0}$ come from Assumption 2. Figure 2 illustrates the modified measurement update (17).
Figure 2: Modified measurement update
It is important to note that the practicality of this pruning algorithm relies on reasonable estimates of $\beta\in\mathcal{KL}$ and $\alpha>0$, which is usually difficult to obtain. For linear systems, they can be obtained easily; however, they are usually quite conservative because of their dependence on the condition number of matrices involved in the estimation. Nonetheless, this technique is useful in providing a theoretical bound of the estimation error.
To this end, we define the estimation error in terms of the center of
$\bar{\mathcal{X}}^{\mathrm{mod}}_{k}:=\langle\hat{c}^{\mathrm{mod}}_{x}(k),\hat{G}^{\mathrm{mod}}_{x}(k)\rangle$ (19)
a zonotope with minimal radius such that it overbounds $\hat{\mathcal{X}}^{\mathrm{mod}}_{k}$ in (17), i.e.,
$\begin{array}[]{rl}\min&\mathrm{rad}(\bar{\mathcal{X}}^{\mathrm{mod}}_{k})\\ \textrm{ subject to }&\hat{\mathcal{X}}^{\mathrm{mod}}_{k}\subseteq\bar{\mathcal{X}}^{\mathrm{mod}}_{k}.\end{array}$ (20)
Since $\lim_{k\rightarrow\infty}\Delta_{k}=\alpha\max(\bar{w},\bar{v})$,
$\lim_{k\rightarrow\infty}\mathrm{rad}(\bar{\mathcal{X}}^{\mathrm{mod}}_{k})\leq\alpha\max(\bar{w},\bar{v}),$
which only depends on the bounds on the process and measurement noise, and not on the attack signals.
6.3 Revised Bound on the Estimation Error
We provide a bound for the estimation error $\hat{c}^{\mathrm{mod}}_{x}(k)-x(k)$, where $\hat{c}^{\mathrm{mod}}_{x}(k)$ is the center of $\bar{\mathcal{X}}^{\mathrm{mod}}_{k}$ in (19).
Proposition 3.
Let Assumptions 1(ii)-(iii), 2, and 3 hold. Consider Algorithm 1 with a modified estimate (17). Then, for every $k\in\mathbb{Z}_{\geq 0}$, the overall estimation error bound is given by the following
$\|\hat{c}^{\mathrm{mod}}_{x}(k)-x(k)\|\leq\Delta_{k}$ (21)
where $\Delta_{k}$ is defined in (18) and $\hat{c}^{\mathrm{mod}}_{x}(k)$ is the center of $\bar{\mathcal{X}}^{\mathrm{mod}}_{k}$ computed by (20).
Proof.
Note that $x(k)\in\hat{\mathcal{X}}_{k}$ according to Theorem 2 and $x(k)\in\mathcal{H}(\hat{x}(k),\Delta_{k})$ according to Assumption 2. Therefore, we have that $x(k)\in\hat{\mathcal{X}}^{\mathrm{mod}}_{k}$. Since $\hat{\mathcal{X}}^{\mathrm{mod}}_{k}\subseteq\bar{\mathcal{X}}^{\mathrm{mod}}_{k}$ because of (20), we obtain (21). ∎
Proposition 3 enables us to conclude that by implementing a secure point-based estimator with properties given by Assumption 2, we can guarantee that the estimation error satisfies (21) with a bound $\Delta_{k}$ that asymptotically converges to a bound that only depends on the bounds on the process and measurement noise, and does not depend on the attack signals $a_{i}(k)$. This is a guarantee that is absent in the purely secure set-based estimation algorithm (Algorithm 1) developed in the previous section. However, we now have a more restrictive Assumption 3 (in comparison to Assumption 1(i)) that guarantees the existence of a secure point-based estimator possessing the property in Assumption 2.
@@ 7 Numerical Simulation
We evaluated our proposed algorithms on two examples. In the first illustrative example, we assume that half of the total number of sensors can be targeted by an attacker at every time instant. We illustrate the estimation algorithm and use a strategy to reduce the complexity with a minimal loss of estimation accuracy. In the second example of a three-story building structure, we assume redundant observability from every subset of two sensors. We use the CORA toolbox [4] to generate our simulations.
(a) Time $k=1$. Sensors 2 and 3 are attacked.
(b) Time $k=2$. Sensors 3 and 4 are attacked.
(c) Time $k=3$. Sensors 1 and 4 are attacked.
Figure 3: Snapshots of estimated sets using Algorithm 1. The attack is generated according to (22) where different sensors are attacked at different time steps.
(a) Time $k=1$. Sensors 2 and 3 are attacked.
(b) Time $k=2$. Sensors 3 and 4 are attacked.
(c) Time $k=3$. Sensors 1 and 4 are attacked.
Figure 4: Snapshots of estimated sets using Algorithm 1 while applying a complexity reduction technique to overapproximate the estimated intersecting sets by one set. The attack is generated according to (22) where different sensors are attacked at different time steps.
@@ 7.1 Two Dimensional Linear System
Let the process noise bound be $\mathcal{W}=\langle 0,\sigma_{\mathcal{W}}I_{2}\rangle$ with $\sigma_{\mathcal{W}}=0.02$ and the measurement noise bounds of the four sensors be $\mathcal{V}_{1}=\mathcal{V}_{2}=\mathcal{V}_{3}=\mathcal{V}_{4}=\langle 0,\sigma_{\mathcal{V}}I_{2}\rangle$ with $\sigma_{\mathcal{V}}=1$, where $I_{2}$ is the identity matrix of dimension $2\times 2$. The attack is generated according to the following relation
@@
(a) Time $k=1$, Sensor 2 attacked.
(b) Time $k=2$, Sensor 3 attacked.
(c) Time $k=3$, Sensor 1 attacked.
Figure 5: Snapshots of estimated sets using Algorithm 1 under time-varying attack, where different sensors are attacked at different time steps.
@@ 7.2 Three-story Building Structure
[…]
@@
We suppose that the attacker can compromise only one sensor at each time, i.e., $q=1$, whose rationale is explained below. Although the pairs $(A,C_{1})$ ,$(A,C_{2})$, and $(A,C_{3})$ are ‘theoretically’ observable, they are not ‘practically’ observable because the corresponding observability matrices have some singular values very close to zero. Therefore, in this example, we assume redundant observability from every pair of two sensors. Our goal is to monitor the building’s floor displacements and velocities irrespective of the compromised sensor.
@@ 8 Conclusion and future work
We presented a secure set-based state estimation algorithm that can ensure the inclusion of the true state of an LTI system in the estimated set even when all but one sensor could be compromised by an adversarial attacker. Our proposed algorithm overcomes the limitation of point-based secure estimators that restrict the number of attacked sensors to strictly less than half of the total number of sensors. We achieved this by constructing agreement sets from the intersection of various observable combinations of measurement update sets. We showed that our algorithm guarantees the inclusion of the true state in the estimated set. Moreover, we proposed sufficient conditions for detecting, identifying, and filtering the attack signals and presented a simple algorithm to identify the set of compromised sensors at every time instant. The proposed algorithm may find applications in safety verification of safety-critical systems whose multiple sensors could be compromised by an attacker.
@@
We also incorporated a secure point-based state estimator into our algorithm to prune the candidate sets when less than half of the sensors can be compromised. This strategy’s effectiveness depends on an accurate approximation of the guaranteed estimation error provided by the secure point-based state estimator. Nonetheless, the modified algorithm provides asymptotic convergence guarantees with an explicit upper bound that depends only the known process and measurement noise, and not on the attack signals.
Our future work will focus on the set-based secure state estimation of nonlinear systems and developing a data-driven approach for secure estimation when the system model is unknown. We also highlight several open questions that remain unaddressed in this paper. For instance, devising better agreement protocols that can improve the lower bounds on the detectable, identifiable, and filterable attack signals will further strengthen the set-based estimation approach presented in this paper. The question of computational complexity under stealthy attacks remains to be addressed in a rigorous manner. The stability of the estimated set is only addressed for the case when strictly less than half of the sensors are attacked. When more than half of the sensors are attacked, it is challenging to upper bound the estimated set by a contracting set. Finally, from the attacker’s perspective, analyzing the complexity of generating stealthy attacks that result in the increased complexity of the set-based estimation is an interesting research problem.

### v3

@@ 5.3 Attack Filtering
In addition to the measurement update, the attacks could also be filtered out via the intersections in the agreement protocol (7), which could yield an empty agreement set $\mathcal{I}_{k}^{h}$ in case of inconsistencies between the measurement updates of the sensors in $\mathsf{J}_{h}$. However, unlike the measurement update, the attacks may not always be identified whenever some of the agreement sets are empty. This is because a compromised sensor $i$ could be a part of multiple agreement sets $\mathcal{I}_{k}^{h}$, for $h\in\mathbb{Z}_{[1,n_{\mathsf{J}}]}$, some of which may turn out to be non-empty. One can identify the attack $a_{i}(k)$ if, for every $h\in\mathbb{Z}_{[1,n_{\mathsf{J}}]}$ such that $i\in\mathsf{J}_{h}$, we have $\mathcal{I}_{k}^{h}=\emptyset$.
@@
$\bar{\mathcal{I}}_{k}^{h,\mathsf{A}_{k}}=\langle c_{\mathcal{I}}^{h,\mathsf{A}_{k}}(k),G_{\mathcal{I}}^{h,\mathsf{A}_{k}}(k)\rangle.$
@@
$\bar{\mathcal{I}}_{k}^{h,\mathsf{S}_{k}}=\langle c_{\mathcal{I}}^{h,\mathsf{S}_{k}}(k),G_{\mathcal{I}}^{h,\mathsf{S}_{k}}(k)\rangle.$
@@
The proof follows from a similar line of argument as the proof of Theorem 4. ∎
@@
Notice that $c_{\mathcal{I}}^{h,\mathsf{A}_{k}}(k)$, which is the center of the overbounding zonotope $\bar{\mathcal{I}}_{k}^{h,\mathsf{A}_{k}}$, is actually the Chebyshev center of the agreement set $\mathcal{I}_{k}^{h,\mathsf{A}_{k}}$. This agreement set is formed by the intersections of the measurement update sets of the attacked sensors $\mathsf{A}_{k}$. The centers of these measurement update sets are influenced by the corresponding attack signals $a_{j}(k)$. Therefore, to achieve $\mathcal{I}_{k}^{h}\neq\emptyset$, it is necessary that the attacker coordinates the attacks on sensors in such a way that the condition of Proposition 5 is violated.
From the attacker’s perspective, the condition of Proposition 5 is necessary for $\mathcal{I}_{k}^{h}\neq\emptyset$, and is not sufficient. This points to several interesting open problems for designing effective attacks. Beyond Theorem 4 and Propositions 3 and 5, our future work will investigate whether a sufficient condition could be derived for the attacker to ensure that both the measurement updates and the agreement sets are non-empty. Moreover, the complexity of generating such attacks is also one of the important questions that will be deferred for our future work.
@@ 5.4 Algorithm for Attack Filtering and Identification
A notable contribution of the set-based state estimation scheme in this paper over other secure schemes is Assumption 1(i), which allows the attacker to compromise not only up to p1p-1 sensors at each time instant but also different subsets of sensors at different times. To this end, we remark that we can detect only those compromised sensors that are injected with non-stealthy attack signals, whereby non-stealthiness we mean the attack signals that satisfy the conditions of Theorem 4 and Propositions 3 and 5. The attack identification algorithm is fairly simple and can be summarized as in Algorithm 2.
@@
The estimated safe subset $\hat{\mathsf{S}}_{k}$ contains the true safe subset $\mathsf{S}_{k}$ at time $k$, i.e., $\mathsf{S}_{k}\subseteq\hat{\mathsf{S}}_{k}$. If there are sensors that are injected with small-valued stealthy attack signals, Algorithm 2 may not identify those attacks and considers those sensors to be uncompromised. Therefore, the estimated set of identified attacked sensors $\hat{\mathsf{A}}_{k}$ is only a subset of the true set of attacked sensors $\mathsf{A}_{k}=\mathbb{Z}_{[1,p]}\setminus\mathsf{S}_{k}$.
@@
In Algorithm 2, we identify a subset of attacked sensors at every time $k$. However, in the time-invariant attack setting where the attacker does not change the subset of sensors to compromise at every time $k$, Algorithm 2 can be adapted to cumulatively detect and remove the attacked sensors over time. This has potential applications in sensor fault detection and isolation, as faults can be considered as naive, time-invariant attacks.
@@ 6 Numerical Simulation
We evaluated our proposed algorithms using two examples. In the first illustrative example, we assume that half of the total number of sensors can be targeted by an attacker at every time instant. We illustrate the estimation algorithm and use a strategy to reduce the complexity with a minimal loss of estimation accuracy. In the second example of a three-story building structure, we assume redundant observability from every subset of two sensors. We use the CORA toolbox [4] to generate our simulations.
@@ 6.1 Two Dimensional Linear System
Let the process noise bound be $\mathcal{W}=\langle 0,\sigma_{\mathcal{W}}I_{2}\rangle$ with $\sigma_{\mathcal{W}}=0.02$ and the measurement noise bounds of the four sensors be $\mathcal{V}_{1}=\mathcal{V}_{2}=\mathcal{V}_{3}=\mathcal{V}_{4}=\langle 0,\sigma_{\mathcal{V}}I_{2}\rangle$ with $\sigma_{\mathcal{V}}=1$, where $I_{2}$ is the identity matrix of dimension $2\times 2$. The attack is generated according to the following
@@
(a) Time $k=1$. Sensors $\{2,3\}$ are attacked.
(b) Time $k=2$. Sensors $\{3,4\}$ are attacked.
(c) Time $k=3$. Sensors $\{1,4\}$ are attacked.
Figure 2: Snapshots of estimated sets using Algorithm 1.
(a) Time $k=1$. Sensors $\{2,3\}$ are attacked.
(b) Time $k=2$. Sensors $\{3,4\}$ are attacked.
(c) Time $k=3$. Sensors $\{1,4\}$ are attacked.
Figure 3: Snapshots of estimated sets using Algorithm 1 while applying a complexity reduction technique to overapproximate the estimated intersecting sets by one set.
@@ 6.2 Three-story Building Structure
$M$ $=\mathrm{diag}([\begin{array}[]{ccc}478350&478350&517790\end{array}])\quad\text{(kg)}$
$D$ $=10^{5}\times\left[\begin{array}[]{ccc}7.7626&-3.7304&0.6514\\ -3.7304&5.8284&-2.0266\\ 0.6514&-2.0266&2.4458\end{array}\right]\quad\text{(Ns/m)}$
$S$ $=10^{8}\times\left[\begin{array}[]{ccc}4.3651&-2.3730&0.4144\\ -2.3730&3.1347&-1.2892\\ 0.4144&-1.2892&0.9358\end{array}\right]\quad\text{(N/m)}$
$G$ $=[\begin{array}[]{ccc}478350&478350&517790\end{array}]^{\top}\quad\text{(kg)}.$
@@
(a) Time $k=1$, Sensor 2 attacked.
(b) Time $k=2$, Sensor 3 attacked.
(c) Time $k=3$, Sensor 1 attacked.
Figure 4: Snapshots of estimated sets using Algorithm 1.
@@
We suppose that the attacker can compromise only one sensor at each time, i.e., $q=1$, whose rationale is explained below. Although the pairs $(A,C_{1})$ $(A,C_{2})$, and $(A,C_{3})$ are ‘theoretically’ observable, they are not ‘practically’ observable because the corresponding observability matrices have some singular values very close to zero. Therefore, in this example, we assume redundant observability from every pair of two sensors. Our goal is to monitor the building’s floor displacements and velocities irrespective of the compromised sensor.
@@ 7 Conclusion and Future Work
We presented a secure set-based state estimation algorithm that can ensure the inclusion of the true state of an LTI system in the estimated set even when all but one sensor could be compromised by an attacker. We achieved this by constructing agreement sets from the intersection of various observable combinations of measurement update sets. We showed that our algorithm guarantees the inclusion of the true state in the estimated set. Moreover, we proposed sufficient conditions for detecting, identifying, and filtering the attack signals and presented a simple algorithm to identify the set of compromised sensors at every time instant. The proposed algorithm may find applications in safety verification of safety-critical systems whose multiple sensors could be compromised by an attacker.
@@
Our future work will focus on the set-based secure state estimation of nonlinear systems and developing a data-driven approach for secure estimation when the system model is unknown. We also highlight several open questions that remain unaddressed in this paper. For instance, devising better agreement protocols that can improve the lower bounds on the detectable, identifiable, and filterable attack signals will further strengthen the set-based estimation approach presented in this paper. The question of computational complexity under stealthy attacks remains to be addressed in a rigorous manner. The stability of the estimated set is only addressed for the case when strictly less than half of the sensors are attacked. When more than half of the sensors are attacked, it is challenging to upper-bound the estimated set by a contracting set. Finally, from the attacker’s perspective, analyzing the complexity of generating stealthy attacks that result in the increased complexity of the set-based estimation is an interesting research problem.
Acknowledgment
This work has received funding from the European Union’s Horizon Research and Innovation Programme under grant agreement No. 830927 and Marie Skłodowska-Curie grant agreement No. 101062523. It is also supported by the Swedish Research Council and the Knut and Alice Wallenberg Foundation, Sweden.